OpenStudy (anonymous):

I need help with a Baseline.. I don't know any of it..<3 @Hero

OpenStudy (anonymous):

At a carnival game, a large horizontal board has holes of the same size cut in it. For the game, you toss a ball onto the board and it lands or rolls into one of the holes. The board is surrounded on all sides, so the ball is certain to fall in a hole. Three of the holes will earn a small prize and one of the holes will earn a large prize. If Grady buys a ticket for one ball that he'll toss on the board, what is the probability that he'll win a prize? A) 5% B)15% C)20% D)40%

OpenStudy (anonymous):

P.s. There a lot more questions I need help with.. -.-

OpenStudy (anonymous):

the way it is written it sounds as if it will surely fall in to one hole and so surely win some prize. i think maybe there is something missing here

OpenStudy (anonymous):

or maybe a number missing here At a carnival game, a large horizontal board has ???? holes of the same size cut in it"

OpenStudy (anonymous):

20 holes.. sorry x)

OpenStudy (anonymous):

Is the answer 5%?

OpenStudy (anonymous):

so if we assume the holes are equally likely to be landed in, it is \(\frac{4}{20}=\frac{20}{100}=20\%\)

OpenStudy (anonymous):

4 holes give prizes, 20 holes total, ratio is \(\frac{4}{20}\)

OpenStudy (anonymous):

How'd you get that answer.. o.o

OpenStudy (anonymous):

Did you mulitply 100 by 20? To get 20 %?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

how did i get what answer?

OpenStudy (anonymous):

i mean, are you asking why it is \(\frac{4}{20}\) or are you asking "why is \(\frac{4}{20}=20\% ?\)"

OpenStudy (anonymous):

No I mean, how did you solve this problem? I totally lost on how to solve it.. how did you get 4/20 and 20/100?

OpenStudy (anonymous):

Also, how did you get that answer? ^

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

how many holes contain prizes?

OpenStudy (anonymous):

..Idk ._.